For the following problem, the slope and one point on a line is given. Use the point-slope form to find the equation of that line. (Write your answer in slope-intercept form.)
(−6,5), m = −4/3
Please help me im so confused :(

Answer by acw1213(28)

"m" is the slope.
We know 4/3 is the slope here.
Point-slope form is
y - y1 = m(x - x1)
When you write point-slope form, the ordered pair they give you is used to plug in for x1 and y1 in point-slope form.
Here, your ordered pair (-6,5)
x1 = -6
y1 = 5
Plug in x1, y1, and the slope into the formula.
y - 5 = 4/3(x - (-6)
Simplify. x - (-6) is the same as x + 6
y - 5 = 4/3(x + 6)
Now we need to get this into slope-intercept form. Distribute 4/3 into the parentheses.
y - 5 = 4/3*x + 4/3*6
Simplify.
4/3*x is the same as 4/3x
4/3*6 is 8.
y - 5 = 4/3x + 8
Slope intercept form is y = mx + b. We need the y to be ALONE on the left side.
Add 5 on both sides. You can only add 5 to its like term, 8.
y = 4/3x + 13
Ta-da! :)